OpenStudy (anonymous):

4x-10y=-3 , 12x=5y=12 What is x? and can u explain plz

OpenStudy (tkhunny):

'x' is a variable or placeholder for an, as yet, unknown quantity. What else would it be?

OpenStudy (anonymous):

idk.

OpenStudy (tkhunny):

Maybe you mean what value for x satisfies the two equations?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

4x + 10y = -3 12x... what is the 2nd equation ? Please write it again

OpenStudy (tkhunny):

4x-10y=-3 12x+5y=12 Your task it to find two values, represented by x and y, that satisfy both equations simultaneously. There are at least a half dozen ways to do this. Which way are you studying? Substitution? Elimination?